# Search for Multiple Adjacent Marked Vertices on the Hypercube by a Quantum Walk with Partial Phase Inversion

### v2

@@ A B C
9 0.981 2 0.981 2 0.948 1
11 0.970 2 0.970 2 0.964 1
Another analysis of the results for the weight $l=n^{2}/N$ was performed. Two scenarios are compared, and some results are shown in Table 2. The results presented in column A refer to the scenario with only non-adjacent marked vertices obtained by Souza et al. (2023). In the scenario presented in column B where we have both types of marked vertices for each $k$ adjacent vertices, we have $k-1$ non-adjacent vertices. Although there are adjacent marked vertices in the sample, the use of multiple self-loops guarantees, in some cases, maximum success probabilities close to $1$.
Note that, in the case where there are only non-adjacent marked vertices, as the number of marked vertices increases, the number of self-loops decreases. However, when there are marked adjacent vertices, a more significant number of self-loops is needed to maintain the success probability close to the maximum. Comparisons made between different scenarios and the same weights show that the type of marked vertex influences the search result. However, although there are adjacent marked vertices in the sample, partial state inversion guarantees, in some cases, maximum success probabilities close to $1$.
Table 2: Comparison between the probability of success and number of self-loops for two different scenarios for weight value $l=n^{2}/N$. Column A represents the results found by Souza et al. (2023) to search for non-adjacent marked vertices, and column B to search for adjacent and non-adjacent marked vertices.
@@
Now, let us analyze the case where the type of marked vertices is the same, but the weights are different. Considering the behavior of the probability of success in Figures 1c and 1d, we can see that not only the type of marked vertices influences the probability of success, but also the weight value. Note that the difference in weight composition, in this case, is the number of marked vertices. We can see that after the increase in the number of self-loops, overall, we significantly improved the probability of success. We can better see these results in Table 3, which shows the maximum probabilities of success and the number of self-loops according to the number of marked vertices and weight value. Comparing the results described in Table 3a and Table 3b, it is essential to realize that the weight composition is very relevant. Although the type of marked vertices can influence the probability of success, with an ideal weight value along with an ideal number of self-loops and partial phase inversion, it is possible to improve the results. The exception was $k=3$, where there was a reduction in the probability of success but still close to $1$. The other bold lines show the cases with the more expressive improvements in the probability of success. In general, there was a significant increase in the number of self-loops.
@@
As in Souza et al. (2023), we analyzed whether the relative position of the non-adjacent marked vertices influences the results of the maximum probability of success. We also used the coefficient of variation to analyze the results’ dispersion level. The relative position of the non-adjacent marked vertices also did not show significant influence considering a numerical precision of four digits. Fig. 2 shows the coefficient of variation for the results presented in Fig. 1. Variations around the mean value are small. However, the behavior exhibited is stable. The maximum success probabilities close to $1$ coincide with these small variations. Considering the weight value of the self-loop, in general, the weight $l=(n^{2}/N)\cdot k$ indicated minor variability.
Figure 2: The coefficient of variation of the MSLQW – PPI to search for adjacent and non-adjacent marked vertices. The results are represented in percentage terms. LABEL:sub@fig:probability-distribution-neighbors-non-neighbors-a-standard-deviation, LABEL:sub@fig:probability-distribution-neighbors-non-neighbors-b-standard-deviation, LABEL:sub@fig:probability-distribution-neighbors-non-neighbors-c-standard-deviation, and LABEL:sub@fig:probability-distribution-neighbors-non-neighbors-d-standard-deviation represents the coefficient of variation of the results presented in Figures 1a, 1b, 1c, and 1d for the weight values $l=n/N$, $l=(n/N)\cdot k$, $l=n^{2}/N$, and $l=(n^{2}/N)\cdot k$, respectively.
@@ 4.2 Analyzing the search with adjacent marked vertices
The simulations performed in the samples of the previous scenario were necessary to obtain the average behavior based on the relative position of the non-adjacent marked vertices. In this scenario, let us analyze only adjacent vertices. According to Nahimovs et al. (2019), a stationary state occurs when two adjacent marked vertices exist. In this case, the maximum probability of success obtained in our simulations was approximately $p=0.02$ for all weights. Now, consider $k\geqslant 3$. Fig. 3a shows the probability of success for the weight $l=n/N$. Comparing the results presented by Souza et al. (2021, 2023) to search for multiple marked vertices and a single self-loop, we had an improvement in the success probability for $k=3$ marked vertices, which were $p=0.745$ and evolved to $p=0.999$ with $m=3$ self-loops.
Fig. 3b shows the probability of success for the weight $l=(n/N)\cdot k$. Compared to the results obtained by Souza et al. (2021) for searching multiple adjacent marked vertices using a single self-loop, there was an improvement in the probability of success. As we can see in Table 4, two results are significant, the search for $k=3$ marked vertices, with $9$ self-loops allowed to increase the probability from $p=0.386$ to $p=0.999$. For $k=4$ marked vertices, with $4$ self-loops permitted to increase the probability from $p=0.639$ to $p=0.996$.
Figure 3: The probability of success of the MSLQW – PPI to search for adjacent marked vertices with $n=12$ and $N=4096$ vertices. LABEL:sub@fig:probability-distribution-neighbors-a weight value $l=n/N$. LABEL:sub@fig:probability-distribution-neighbors-b weight value $l=(n/N)\cdot k$. LABEL:sub@fig:probability-distribution-neighbors-c weight value $l=n^{2}/N$. LABEL:sub@fig:probability-distribution-neighbors-d weight value $l=(n^{2}/N)\cdot k$.
Again, let us analyze two different scenarios for the same weight values. Fig. 3c shows the success probabilities for searching only adjacent marked vertices while Fig. 1c shows the success probabilities for searching adjacent and non-adjacent marked vertices both using the weight $l=n^{2}/N$. Note that the behaviors are very similar. However, in the scenario where there are only adjacent marked vertices, which is the case in Fig. 3c, a more significant number of self-loops is necessary when the density of adjacent marked vertices is small. Table 5 compares the success probabilities and the number of self-loops for the cases where the number of marked vertices is the same. Again, note that the results are similar except for $k=3$, where there was a significant increase in the number of self-loops.
@@
Now, consider Fig. 3d. It shows the success probabilities to search for adjacent marked vertices using the self-loop weight $l=(n^{2}/N)\cdot k$. Compared with the results of the scenario presented in Fig. 1d, we notice a very similar behavior where for a small density of marked vertices, a greater number of self-loops is necessary. However, when this density of marked vertices increases, the number of self-loops decreases to the point of approaching the results presented by Souza et al. (2023) for the same weight value $l=(n^{2}/N)\cdot k$. Table 6 shows the number of self-loops needed to obtain the maximum probabilities of success. Comparing with the results presented in Table 3b for the same numbers of marked vertices, it is possible to see that, a greater number of self-loops are required to achieve success probabilities close to $1$ when there are only adjacent marked vertices. However, for $k=3$, $m=30$ was insufficient.
Table 5: Comparison between the probability of success and number of self-loops for two scenarios. The column for 1c represents the results of searching for adjacent and non-adjacent vertices, and the column for 3c to search for adjacent vertices.
@@ 5 Conclusions
In this work, we analyzed the application of MSLQW – PPI in two scenarios based on the type of marked vertices: adjacent and non-adjacent. In the first scenario, the two kinds of marked vertices were searched. Here, we analyzed the relative position of the non-adjacent marked vertices to verify your influence on the results about adjacent vertices. For this, the coefficient of variation was also used to verify the dispersion of results around the maximum success probability mean value as Souza et al. (2023). In the second scenario, we analyzed only the adjacent marked vertices. The dependence on the self-loop weight value is inherent to the lackadaisical quantum walk. Therefore, the composition of the weights is essential. Thus, all analyses were made considering the four weight values. However, when applied to MSLQW – PPI, a strategy of weight distribution is equally necessary because of the multiple self-loop. The strategy of weight distribution in this work was the same used by Souza et al. (2023), i.e., the weight $l=l^{\prime}/m$.
@@
In the first scenario, to search for adjacent and non-adjacent vertices, according to the results, the relative position of non-adjacent marked vertices does not have a significant influence, considering a numerical precision of four digits. The results obtained by Souza et al. (2021, 2023) to search for non-adjacent marked vertices indicate that the weight values influenced the maximum success probabilities according to weight values $l=(n/N)\cdot k$ and $l=(n^{2}/N)\cdot k$. Moreover, when we analyzed the coefficient of variation, we saw that a minor variability coincides with the maximum probability of success. Fig. 2 shows that the number of marked vertices influences the results causing a more significant variability. In the second scenario, the results obtained in the search for adjacent marked vertices present very similar results to the search for adjacent and non-adjacent marked vertices, except in some cases. For example, when the weight $l=(n^{2}/N)\cdot k$, the number of self-loops increases considerably.
@@
In summary, we conclude that for MSLQW – PPI, there is a dependence between the weight value, the vertex type, the number of marked vertices, and the number of self-loops needed to obtain success probabilities close to $1$. In the search for adjacent and non-adjacent marked vertices, the weight that presented the best results was $l=(n^{2}/N)\cdot k$. In the search for adjacent marked vertices, three of four weights presented the best results: $l=(n/N)\cdot k$, $l=(n^{2}/N)$, and $l=(n^{2}/N)\cdot k$. The results presented in Fig. 3d show that from a certain $k$, the self-loops converge to a certain quantity. In future works, we intend to apply this methodology to evaluate the MSLQW – PPI in other $d$-regular structures with samples that contain adjacent marked vertices. We intend to verify the convergence of the number of multiple self-loops for a specific $m$ from a certain $k$ for the weight value $l=(n^{2}/N)\cdot k$.

### v3

@@ A B C
9 - - 0.981 2 0.948 1
11 - - 0.970 2 0.964 1
Fig. 1c shows the probability of success in the search for adjacent and non-adjacent vertices using the weight value $l=n^{2}/N$. This weight is proposed by Souza et al. (2023) and is composed of the weight value proposed by Rhodes and Wong (2020) to search for one marked vertex plus an exponent in the element that represents the degree of the vertex in the numerator. Compared with the results found by Souza et al. (2021) and with the results shown in Fig. 1a, there was a significant improvement in the maximum probability of success for numbers $k>3$ marked vertices. In this scenario, the success probabilities depend on the inversely proportional relationship between the number $k$ of marked vertices and the number $m$ of Self-loops. This means that as the number of marked vertices increases, the number of self-loops decreases and the other way around, however, the maximum probability of success continues above $p=0.97$.
Another analysis of the results for the weight $l=n^{2}/N$ was performed. Two different scenarios are compared and some results are shown in Table 4. The results presented in column A refer to the scenario with only non-adjacent marked vertices obtained by Souza et al. (2023). In the scenario presented in column B where we have both types of marked vertices for each $k$ adjacent vertices, we have $k-1$ non-adjacent vertices. Although there are adjacent marked vertices in the sample, the use of multiple self-loops guarantees, in some cases, maximum success probabilities close to $1$.
Note that, in the case where there are only non-adjacent marked vertices, as the number of marked vertices increases the number of self-loops decreases. However, when there are marked adjacent vertices, a larger number of self-loops is needed to maintain the success probability close to the maximum. Comparisons made between different scenarios and the same weights show that the type of marked vertex influences the search result. However, although there are adjacent marked vertices in the sample, partial state inversion guarantees, in some cases, maximum success probabilities close to $1$.
Table 4: Comparison between the probability of success and number of self-loops for two different scenarios for weight value $l=n^{2}/N$. Column A represents the results found by Souza et al. (2023) to search for non-adjacent marked vertices and column B to search for adjacent and non-adjacent marked vertices.
@@
Now, we are going to analyze the case where the scene is the same but the weights are different. Considering the behavior of the probability of success in Figures 1c and 1d, we can see that not only the type of marked vertices influences the probability of success, but also the weight value. Note that the difference in weight composition, in this case, is the number of marked vertices. We can see that after the increase in the number of self-loops, overall we had a significant improvement in the probability of success. We can better see these results in Table 5 with shows the maximum probabilities of success and the number of self-loops according to the number of marked vertices and weight value. Comparing the results described in Table 5a and Table 5b, it is important to realize that the weight composition is very relevant. Although the type of marked vertices can influence the probability of success, with an ideal weight value along with an ideal number of self-loops, it is possible to improve the results. The exception was $k=3$, where there was a reduction in the probability of success but still close to $1$. The other bold lines show the cases with the more expressive improvements in the probability of success. In general, there was a significant increase in the number of self-loops.
@@
As in Souza et al. (2023), we analyzed whether the relative position of the non-adjacent marked vertices influences the results of the maximum probability of success. We also used the coefficient of variation to analyze the level of dispersion of the results. The relative position of the non-adjacent marked vertices also did not show significant influence considering a numerical precision of four digits. Fig. 2 shows the coefficient of variation for the results presented in Fig. 1. Variations around the mean value are small, however, the behavior shown is stable. The maximum success probabilities close to $1$ coincide with these small variations. Considering the weight value of the self-loop, in general, the weight $l=(n^{2}/N)\cdot k$ indicated minor variability.
(a) $l=n/N$
(b) $l=(n/N)\cdot k$
(c) $l=n^{2}/N$
(d) $l=(n^{2}/N)\cdot k$
Figure 2: The coefficient of variation of the MSLQW-PPI to search for adjacent and non-adjacent marked vertices. The results are represented in percentage terms. LABEL:sub@fig:probability-distribution-neighbors-non-neighbors-a-standard-deviation, LABEL:sub@fig:probability-distribution-neighbors-non-neighbors-b-standard-deviation, LABEL:sub@fig:probability-distribution-neighbors-non-neighbors-c-standard-deviation, and LABEL:sub@fig:probability-distribution-neighbors-non-neighbors-d-standard-deviation represents the coefficient of variation of the results presented in Figures 1a, 1b, 1c, and 1d for the weight values $l=n/N$, $l=(n/N)\cdot k$, $l=n^{2}/N$, and $l=(n^{2}/N)\cdot k$, respectively.
@@ 4.2 Analyzing the search with adjacent marked vertices
The simulations performed in the samples of the previous scenario were necessary so that we could obtain the average behavior based on the relative position of the non-adjacent marked vertices. In this scenario, let us analyze only adjacent vertices. According to Nahimovs et al. (2019), when there are two adjacent marked vertices, a stationary state occurs. In this case, the maximum probability of success obtained in our simulations was approximately $p=0.02$ for all weights. Now, consider $k\geqslant 3$. Fig. 3a shows the probability of success for the weight $l=n/N$. Comparing the results presented by Souza et al. (2021) and Souza et al. (2023) to search for multiple marked vertices and a single self-loop we had an improvement in the success probability for $k=3$ marked vertices which were $p=0.745$ and evolved to $p=0.999$ with $m=3$ self-loops.
Fig. 3b shows the probability of success for the weight $l=(n/N)\cdot k$. Compared to the results obtained by Souza et al. (2021) for searching multiple adjacent marked vertices using a single self-loop, there was an improvement in the probability of success. As we can see in Table 6, two results are significant, the search for $k=3$ marked vertices, with $9$ self-loops allowed to increase the probability from $p=0.386$ to $p=0.999$. For $k=4$ marked vertices, with $4$ self-loops allowed to increase the probability from $p=0.639$ to $p=0.996$. Fig. 1b shows the results where both adjacent and non-adjacent vertices are marked. Analyzing the cases where the number of marked vertices is the same, i.e., $k=\{3,5,7,9,11\}$ with $m=2$ self-loops, although the scenarios are different, the maximum probability of success remained above $p=0.99$.
(a) $l=n/N$
(b) $l=(n/N)\cdot k$
(c) $l=n^{2}/N$
(d) $l=(n^{2}/N)\cdot k$
Figure 3: The probability of success of the MSLQW-PPI to search for adjacent marked vertices with $n=12$ and $N=4096$ vertices. LABEL:sub@fig:probability-distribution-neighbors-a weight value $l=n/N$. LABEL:sub@fig:probability-distribution-neighbors-b weight value $l=(n/N)\cdot k$. LABEL:sub@fig:probability-distribution-neighbors-c weight value $l=n^{2}/N$. LABEL:sub@fig:probability-distribution-neighbors-d weight value $l=(n^{2}/N)\cdot k$.
Again, let us analyze two different scenarios for the same weight values. Fig. 3c shows the success probabilities for searching only adjacent marked vertices while Fig. 1c shows the success probabilities for searching adjacent and non-adjacent marked vertices both using the weight $l=n^{2}/N$. Note that the behaviors are very similar, however, in the scenario where there are only adjacent marked vertices, which is the case in Fig. 3c, a greater number of self-loops is necessary when the density of adjacent marked vertices is small. Table 7 shows the comparison between the success probabilities and the number of self-loops for the cases where the number of marked vertices is the same. Again, note that the results are similar except for $k=3$, where there was a significant increase in the number of self-loops.
@@
Now, consider Fig. 3d. It shows the success probabilities to search for adjacent marked vertices using the self-loop weight $l=(n^{2}/N)\cdot k$. Compared with the results of the scenario presented in Fig. 1d we notice a very similar behavior where for a small density of marked vertices a greater number of self-loops is necessary, however, when this density of marked vertices increases, the number of self-loops decreases to the point of approaching the results presented by Souza et al. (2023) for the same weight value $l=(n^{2}/N)\cdot k$. Table 8 shows the number of self-loops needed to obtain the maximum probabilities of success. Comparing with the results presented in Table 5b for the same numbers of marked vertices, it is possible to see that, a greater number of self-loops are required to achieve success probabilities close to $1$ when there are only adjacent marked vertices. However, for $k=3$, $m=30$ was insufficient.
To obtain the complexity of the algorithm proposed by Souza et al. (2023). applied to the search for adjacent vertices, two analyses were performed. The first analysis described in Fig. 4 shows the runtime complexity when $N=2^{n}$ is changed. The second analysis described in Fig. 5 shows how the runtime complexity behaves when $m$ self-loops are added at each vertex of the hypercube. As in Souza et al. (2023), the computational cost when we consider the hypercube size is $O(\sqrt{((n+m)})$. The computational cost where $m$ self-loops varies is $O(\log{(m)})$.
(a) Adjacents.
(b) Adjacents and non adjacents.
Figure 4: The time complexity of the algorithm relative to the size of the hypercube. The solid red line represents the estimated curve and the blue dots are the numerical simulation values of the quantum walk.
(a) Adjacents
(b) Adjacents and non adjacents
(c) Adjacents
(d) Adjacents and non adjacents
Figure 5: The algorithm’s time complexity concerning the number of self-loops at each vertex. The solid lines represent the estimated curves. The points are the values from numerically simulating the quantum walk. For each figure, $n$ is a constant.
Table 7: Comparison between the probability of success and number of self-loops for two different scenarios. The column for 1c represents the results of searching for adjacent and non-adjacent vertices and the column for 3c to search for adjacent vertices.
@@ 5 Conclusions
In this work, we analyzed the application of MSLQW-PPI in two scenarios based on the type of marked vertices: adjacent and non-adjacent. In the first scenario, the two types of marked vertices were searched. Here, we analyzed the relative position of the non-adjacent marked vertices to verify your influence on the results about adjacent vertices, for this, the coefficient of variation was also used to verify the dispersion of results around the maximum success probability mean value as Souza et al. (2023). In the second scenario, we analyzed only the adjacent marked vertices. The dependence on the self-loop weight value is inherent to the lackadaisical quantum walk. Therefore, the composition of the weights is important. Thus, all analyses were made considering the four weight values. However, when applied to MSLQW-PPI a strategy of weight distribution is equally necessary because due to the use of the multiple self-loop. The strategy of weight distribution in this work was the same used by Souza et al. (2023), i.e., the weight $l=l^{\prime}/m$.
@@
In the first scenario, to search for adjacent and non-adjacent vertices, according to the results, the relative position of non-adjacent marked vertices does not have a significant influence, considering a numerical precision of four digits. The results obtained by Souza et al. (2021, 2023) to search for non-adjacent marked vertices indicate that the weight values influenced the maximum success probabilities according to weight values $l=(n/N)\cdot k$ and $l=(n^{2}/N)\cdot k$. Moreover, when we analyze the coefficient of variation, we saw that a minor variability coincides with the maximum probability of success. Fig. 2 shows that the number of marked vertices influences the results causing a more significant variability. In the second scenario, the results obtained in the search for adjacent marked vertices present very similar results to the search for adjacent and non-adjacent marked vertices, except in some cases. For example, when the weight $l=(n^{2}/N)\cdot k$, the number of self-loops increases considerably.
@@
In summary, we conclude that for MSLQW-PPI, there is a dependence between the weight value, the vertex type, the number of marked vertices, and the number of self-loops needed to obtain success probabilities close to $1$. In the search for adjacent and non-adjacent marked vertices, the weight that presented the best results was $l=(n^{2}/N)\cdot k$. In the search for adjacent marked vertices, three of four weights presented the best results: $l=(n/N)\cdot k$, $l=(n^{2}/N)$, and $l=(n^{2}/N)\cdot k$. The results presented in Fig. 3d show that from a certain $k$, the self-loops converge to a certain quantity. In future works, we intend to apply this methodology to evaluate the MSLQW-PPI in other $d$-regular structures with samples that contain adjacent marked vertices. We intend to verify the convergence of the number of multiple self-loops for a specific $m$ from a certain $k$ for the weight value $l=(n^{2}/N)\cdot k$ to search for adjacents marked vertices.
Acknowledgments
Acknowledgments to the Science and Technology Support Foundation of Pernambuco (FACEPE) Brazil, Brazilian National Council for Scientific and Technological Development (CNPq), and Coordenação de Aperfeiçoamento de Pessoal de Nível Superior - Brasil (CAPES) - Finance Code 001 by their financial support to the development of this research.